### v1

@@ 1. Introduction
Large language models (LLMs) have demonstrated extraordinary versatility in various natural language processing (NLP) tasks, subsuming pipelines that were originally tailormade for each task. Despite being trained on massive text corpora, LLMs still face memory-related challenges such as out-of-date and out-of-domain knowledge, and they occasionally hallucinate non-factual or non-sensical content (Zhou et al., 2021; Maynez et al., 2020). To enhance the accuracy and reliability of LLM-generated outputs, retrieval-augmented generation (RAG) has emerged as a promising solution for knowledge-intensive tasks (Zhu et al., 2023; Gao et al., 2023b; Li et al., 2022) (e.glet@tokeneonedotopen-domain question and answering). RAG systems typically follow a “retrieve-then-generate” paradigm (Shao et al., 2023), where a retriever identifies relevant information from an external corpus and uses this information to augment context in constituting the input to a generative model (i.elet@tokeneonedot, generator), thus yielding an improved answer.
@@ 2.6. Reinforcement Learning Alignment
Reward estimation. The downstream generator $\mathcal{G}$ serves as a reward model, the generation of Oreo by policy $\pi_{\_}{\theta_{\_}{t}}$ is passed to $\mathcal{G}$ with query $q$ to generate the answer $y$. When the end of sentence (e.glet@tokeneonedot¡EOS¿) token is generated, the corresponding reward $R_{\_}{t}$ is obtained by comparing the generated answer $y$ with ground truth answer $y_{\_}{gold}$, which is measured by the ROUGE score $R_{\_}{t}=ROUGE(y,y_{\_}{gold})$. However, $\mathcal{G}$ generates answers only after completing all tokens, but (3) updates every action step. To address this, we incorporate a token-level weighting mechanism (Yang et al., 2023). Considering that a token $t$ with higher generation probability deemed more critical by the current policy. Consequently, the token’s contribution to the final reward is proportionally adjusted. We estimate the reward at each step $t$ using the formulation:
@@ 4.1. Effectiveness Evaluation
Regarding extraction and compression baselines, the variation in performance reflects the differing strengths and limitations of these methods in addressing specific tasks. As reported in Table 4, the CXMI-guided model demonstrates performance closest to Oreo across most datasets among all baselines. In contrast, Selective-Context exhibits the lowest performance. This underperformance likely arises from its reliance on the self-information of lexical units (e.glet@tokeneonedottokens, phrases, or sentences), which overlooks the interdependence between compressed content and fails to create a cohesive context for downstream tasks. LLMLingua, which is based on perplexity, also faces challenges due to its lack of contextual modeling, as it does not explicitly consider contextual information during content selection. Furthermore, both LLMLingua and Selective-Context are not specifically designed for RAG scenarios. As a result, these methods lack query awareness, which limits their ability to tailor extracted context to the unique demands of QA tasks with RAG.
@@ 4.2. Efficiency Evaluation
We evaluate the efficiency of Oreo from two aspects: the reduction in the input context length provided to the generator and its impact on QA task performance. Figure 5 provides an overview of the number of tokens sent to the downstream generator when using query-only inputs, full content,and context reconstructed by Oreo . Oreo achieves a remarkable 84% to 94% reduction in token counts while delivering a performance improvement ranging from 8.76% to 37.46% compared to using the original full content. The gains are particularly notable in extractive QA tasks (e.glet@tokeneonedotNQ, TriviaQA) as shown from the steeper improvement in Figure 5 from right endpoints to peaks. The high compression rate and improved performance demonstrates Oreo ’s capability to effectively condense the retrieved context by preserving only the most critical evidence required for accurate answer generation. This also indicates the context reconstructed by Oreo is highly utilized by the generator.
@@ 4.4. Generalizability Evaluation
To evaluate the cross-dataset generalizability of Oreo , we assess its transferring capability by applying models trained on one dataset to tasks in a different dataset without any fine-tuning. This approach tests Oreo ’s ability to generalize its context reconstruction and synthesis capabilities to unseen query distributions. Specifically, we examine performance when using a model trained on PopQA to generate answers for NQ and a model trained on 2QWA to process HQA queries. The results, presented in Table 6, reveal that Oreo achieves competitive performance in zero-shot setting. For example, the model trained on PopQA achieves a score of 0.4352 when applied to NQ, only slightly lower than the performance of being specific trained (i.elet@tokeneonedot0.4413 and 0.4682). Similarly, using the 2WQA-trained model on HQA yields a score of 0.6344, closely matching the intra-dataset score of 0.6384. These findings demonstrate Oreo ’s ability to generalize its context reconstruction to similar QA types effectively, even under query distribution shifts. Its strong performance across datasets highlights its robustness and adaptability, making it a promising solution for open-domain QA tasks that require flexibility in handling diverse knowledge sources and query structures.
@@ 5.1. Post-retrieval Enhancement for RAG
Reranking. Rerankers reorder and prioritize retrieved documents to emphasize the most relevant results. They typically operate sequentially or iteratively after retrieval, leveraging various criteria such as semantic relevance between query and passages (Glass et al., 2022; Hofstätter et al., 2023), connections among documents (Dong et al., 2024), the majority of reader predictions (Mao et al., 2021), and utility for generation (Ma et al., 2023a). Rerankers are usually based on cross-encoder (e.glet@tokeneonedotBGE (Xiao et al., 2023), Mixedbread (Li and Li, 2023)), multi-vector models (e.glet@tokeneonedotColBert (Khattab and Zaharia, 2020; Santhanam et al., 2022)). Recent works explore using LMs as rerankers (e.glet@tokeneonedot, RankT5 (Zhuang et al., 2023), RankZephyr (Pradeep et al., 2023), RankGPT (Sun et al., 2023b)).
@@ 5.2. Reinforcement Learning for Large Language Models
Reinforcement learning for Language Models (RL4LM) has emerged as a transformative technique to further enhance LLMs’ performance during the post-training process (Cao et al., 2024b; Pternea et al., 2024). Traditional RL4LM usually involves a reward model, for example using PPO (Schulman et al., 2015) to update the policy model (e.glet@tokeneonedotInstructGPT (Ouyang et al., 2022), GPT-4 (Achiam et al., 2023)). Some RL4LM such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Reward-aware Preference Optimization (RPO) (Adler et al., 2024) get rid of reward model to provide more stable and computationally efficient solutions (e.glet@tokeneonedotQwen 2 (Chu et al., 2024) and Nemotron-4 (Adler et al., 2024)). Common goals of RL4LM include improving performance of downstream NLP tasks (Deng et al., 2022; Ghalandari et al., 2022; Ramamurthy et al., 2023), minimizing data and resource dependencies (Zhang et al., 2022), aligning model outputs with user intent, values and goals (Ouyang et al., 2022), and adhering to responsible AI principles (Bai et al., 2022a, b). Human feedback can be integrated into the framework by constructing preference datasets, which are then used to fine-tune both the policy and reward models (also termed as Reinforcement Learning from Human Feedback (RLHF)) (Bai et al., 2022a; Ouyang et al., 2022; Hu et al., 2023). Some studies also explore RL4LM without human feedback (Rafailov et al., 2024) or replaced with AI feedback (Bai et al., 2022b; Yuan et al., 2024) by distillation from LLMs (Cui et al., 2023; Park et al., 2024), prompting LLMs as reward functions (Kwon et al., 2023; Lee et al., 2023; Zhang et al., 2024a), and self-rewarding (Yuan et al., 2024), or using performance-based metrics such as fluency or coherence (Ghalandari et al., 2022), and task-specific constraints over the distribution of language (Ramamurthy et al., 2023; Wu et al., 2024b). In the specific domain of RAG, RRAML (Bacciu et al., 2023) employs RL to train an efficient retriever to search for relevant information in in arbitrarily large databases. PRCA (Yang et al., 2023) applies RL to further fine-tune the context extracted by a model pre-trained on domain-specific abstractive summarization tasks to optimize the reward for the generator. BIDER (Jin et al., 2024) adopts RL to bridge the inconsistency between the retriever and generator.
@@ 6. Conclusion
Limitations and future works. Despite Oreo ’s demonstrated effectiveness in short-form factual QA tasks, certain limitations remain. First, to achieve a high compression rate, Oreo enforces a maximum token count during context synthesis. While effective for most factual QA tasks, this constraint may inadvertently exclude vital information, particularly in complex tasks that require multi-hop reasoning or synthesizing extensive, interconnected evidence. Future work should investigate the trade-off between token counts and different tasks, and explore dynamic token allocation strategies that adapt to the query type (e.glet@tokeneonedotmulti-hop QA, long-form QA) and consider latency requirements during inference. Second, while Oreo has shown resilience to perturbed document orders and noise-injected context, it has not been thoroughly tested in adversarial environments, such as those involving misinformation or conflicting evidence in retrieved documents. Future research should consider evaluating and improving Oreo ’s robustness in such challenging scenarios. Third, the current evaluation of Oreo relies heavily on the performance of downstream tasks, which serve as indirect assessments and may introduce bias. Developing advanced evaluation schemes (e.glet@tokeneonedotLLM-as-a-jedge (Gu et al., 2024)), will be essential for providing a more direct and reliable assessment of the quality of Oreo ’s reconstructions. Addressing these challenges could unlock the full potential of Oreo in more diverse and demanding real-world applications.

### v2

@@ 1. Introduction
Large language models (LLMs) have demonstrated extraordinary versatility in various natural language processing (NLP) tasks, subsuming pipelines that were originally tailormade for each task. Despite being trained on massive text corpora, LLMs still face memory-related challenges such as out-of-date and out-of-domain knowledge, and they occasionally hallucinate non-factual or non-sensical content (Zhou et al., 2021; Maynez et al., 2020). To enhance the accuracy and reliability of LLM-generated outputs, retrieval-augmented generation (RAG) has emerged as a promising solution for knowledge-intensive tasks (Zhu et al., 2023; Gao et al., 2023b; Li et al., 2022) (e.g open-domain question and answering). RAG systems typically follow a “retrieve-then-generate” paradigm (Shao et al., 2023), where a retriever identifies relevant information from an external corpus and uses this information to augment context in constituting the input to a generative model (i.e, generator), thus yielding an improved answer.
@@ 2.6. Reinforcement Learning Alignment
Reward estimation. The downstream generator $\mathcal{G}$ serves as a reward model, the generation of Oreo by policy $\pi_{\_}{\theta_{\_}{t}}$ is passed to $\mathcal{G}$ with query $q$ to generate the answer $y$. When the end of sentence (e.g ¡EOS¿) token is generated, the corresponding reward $R_{\_}{t}$ is obtained by comparing the generated answer $y$ with ground truth answer $y_{\_}{gold}$, which is measured by the ROUGE score $R_{\_}{t}=ROUGE(y,y_{\_}{gold})$. However, $\mathcal{G}$ generates answers only after completing all tokens, but (3) updates every action step. To address this, we incorporate a token-level weighting mechanism (Yang et al., 2023). Considering that a token $t$ with higher generation probability deemed more critical by the current policy. Consequently, the token’s contribution to the final reward is proportionally adjusted. We estimate the reward at each step $t$ using the formulation:
@@ 4.1. Effectiveness Evaluation
Regarding extraction and compression baselines, the variation in performance reflects the differing strengths and limitations of these methods in addressing specific tasks. As reported in Table 4, the CXMI-guided model demonstrates performance closest to Oreo across most datasets among all baselines. In contrast, Selective-Context exhibits the lowest performance. This underperformance likely arises from its reliance on the self-information of lexical units (e.g tokens, phrases, or sentences), which overlooks the interdependence between compressed content and fails to create a cohesive context for downstream tasks. LLMLingua, which is based on perplexity, also faces challenges due to its lack of contextual modeling, as it does not explicitly consider contextual information during content selection. Furthermore, both LLMLingua and Selective-Context are not specifically designed for RAG scenarios. As a result, these methods lack query awareness, which limits their ability to tailor extracted context to the unique demands of QA tasks with RAG.
@@ 4.2. Efficiency Evaluation
We evaluate the efficiency of Oreo from two aspects: the reduction in the input context length provided to the generator and its impact on QA task performance. Figure 5 provides an overview of the number of tokens sent to the downstream generator when using query-only inputs, full content,and context reconstructed by Oreo . Oreo achieves a remarkable 84% to 94% reduction in token counts while delivering a performance improvement ranging from 8.76% to 37.46% compared to using the original full content. The gains are particularly notable in extractive QA tasks (e.g NQ, TriviaQA) as shown from the steeper improvement in Figure 5 from right endpoints to peaks. The high compression rate and improved performance demonstrates Oreo ’s capability to effectively condense the retrieved context by preserving only the most critical evidence required for accurate answer generation. This also indicates the context reconstructed by Oreo is highly utilized by the generator.
@@ 4.4. Generalizability Evaluation
To evaluate the cross-dataset generalizability of Oreo , we assess its transferring capability by applying models trained on one dataset to tasks in a different dataset without any fine-tuning. This approach tests Oreo ’s ability to generalize its context reconstruction and synthesis capabilities to unseen query distributions. Specifically, we examine performance when using a model trained on PopQA to generate answers for NQ and a model trained on 2QWA to process HQA queries. The results, presented in Table 6, reveal that Oreo achieves competitive performance in zero-shot setting. For example, the model trained on PopQA achieves a score of 0.4352 when applied to NQ, only slightly lower than the performance of being specific trained (i.e 0.4413 and 0.4682). Similarly, using the 2WQA-trained model on HQA yields a score of 0.6344, closely matching the intra-dataset score of 0.6384. These findings demonstrate Oreo ’s ability to generalize its context reconstruction to similar QA types effectively, even under query distribution shifts. Its strong performance across datasets highlights its robustness and adaptability, making it a promising solution for open-domain QA tasks that require flexibility in handling diverse knowledge sources and query structures.
@@ 5.1. Post-retrieval Enhancement for RAG
Reranking. Rerankers reorder and prioritize retrieved documents to emphasize the most relevant results. They typically operate sequentially or iteratively after retrieval, leveraging various criteria such as semantic relevance between query and passages (Glass et al., 2022; Hofstätter et al., 2023), connections among documents (Dong et al., 2024), the majority of reader predictions (Mao et al., 2021), and utility for generation (Ma et al., 2023a). Rerankers are usually based on cross-encoder (e.g BGE (Xiao et al., 2023), Mixedbread (Li and Li, 2023)), multi-vector models (e.g ColBert (Khattab and Zaharia, 2020; Santhanam et al., 2022)). Recent works explore using LMs as rerankers (e.g, RankT5 (Zhuang et al., 2023), RankZephyr (Pradeep et al., 2023), RankGPT (Sun et al., 2023b)).
@@ 5.2. Reinforcement Learning for Large Language Models
Reinforcement learning for Language Models (RL4LM) has emerged as a transformative technique to further enhance LLMs’ performance during the post-training process (Cao et al., 2024b; Pternea et al., 2024). Traditional RL4LM usually involves a reward model, for example using PPO (Schulman et al., 2015) to update the policy model (e.g InstructGPT (Ouyang et al., 2022), GPT-4 (Achiam et al., 2023)). Some RL4LM such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Reward-aware Preference Optimization (RPO) (Adler et al., 2024) get rid of reward model to provide more stable and computationally efficient solutions (e.g Qwen 2 (Chu et al., 2024) and Nemotron-4 (Adler et al., 2024)). Common goals of RL4LM include improving performance of downstream NLP tasks (Deng et al., 2022; Ghalandari et al., 2022; Ramamurthy et al., 2023), minimizing data and resource dependencies (Zhang et al., 2022), aligning model outputs with user intent, values and goals (Ouyang et al., 2022), and adhering to responsible AI principles (Bai et al., 2022a, b). Human feedback can be integrated into the framework by constructing preference datasets, which are then used to fine-tune both the policy and reward models (also termed as Reinforcement Learning from Human Feedback (RLHF)) (Bai et al., 2022a; Ouyang et al., 2022; Hu et al., 2023). Some studies also explore RL4LM without human feedback (Rafailov et al., 2024) or replaced with AI feedback (Bai et al., 2022b; Yuan et al., 2024) by distillation from LLMs (Cui et al., 2023; Park et al., 2024), prompting LLMs as reward functions (Kwon et al., 2023; Lee et al., 2023; Zhang et al., 2024a), and self-rewarding (Yuan et al., 2024), or using performance-based metrics such as fluency or coherence (Ghalandari et al., 2022), and task-specific constraints over the distribution of language (Ramamurthy et al., 2023; Wu et al., 2024b). In the specific domain of RAG, RRAML (Bacciu et al., 2023) employs RL to train an efficient retriever to search for relevant information in in arbitrarily large databases. PRCA (Yang et al., 2023) applies RL to further fine-tune the context extracted by a model pre-trained on domain-specific abstractive summarization tasks to optimize the reward for the generator. BIDER (Jin et al., 2024) adopts RL to bridge the inconsistency between the retriever and generator.
@@ 6. Conclusion
Limitations and future works. Despite Oreo ’s demonstrated effectiveness in short-form factual QA tasks, certain limitations remain. First, to achieve a high compression rate, Oreo enforces a maximum token count during context synthesis. While effective for most factual QA tasks, this constraint may inadvertently exclude vital information, particularly in complex tasks that require multi-hop reasoning or synthesizing extensive, interconnected evidence. Future work should investigate the trade-off between token counts and different tasks, and explore dynamic token allocation strategies that adapt to the query type (e.g multi-hop QA, long-form QA) and consider latency requirements during inference. Second, while Oreo has shown resilience to perturbed document orders and noise-injected context, it has not been thoroughly tested in adversarial environments, such as those involving misinformation or conflicting evidence in retrieved documents. Future research should consider evaluating and improving Oreo ’s robustness in such challenging scenarios. Third, the current evaluation of Oreo relies heavily on the performance of downstream tasks, which serve as indirect assessments and may introduce bias. Developing advanced evaluation schemes (e.g LLM-as-a-jedge (Gu et al., 2024)), will be essential for providing a more direct and reliable assessment of the quality of Oreo ’s reconstructions. Addressing these challenges could unlock the full potential of Oreo in more diverse and demanding real-world applications.